### v1

# Matter-Mediated Entanglement in Classical Gravity: Suppression by Binding Potentials and Localization

Abstract
Aziz and Howl [Nature 646, 813–817 (2025)] argue that two spatially separated masses can become entangled even when gravity is treated as a classical field, by invoking higher-order “virtual-matter” processes in a QFT description of matter, which is non-LOCC (local operations and classical communication). We point out that the relevant mechanism is not intrinsically field-theoretic, but is essentially a quantum tunneling/evanescent matter channel, which is already captured within ordinary quantum mechanics. More importantly, the microscopic constituents of realistic macroscopic objects are bound and localized by strong potentials, introducing a large internal energy scale that suppresses coherent propagation between distant bodies. Including such binding/localization generically yields an exponential suppression, rendering the matter-mediated contribution negligible at the macroscopic separations relevant to gravitational-entanglement proposals. Consequently, the entanglement identified by AH diagnoses the presence of a coherent matter-exchange channel rather than the classical or quantum nature of gravity, and it does not undermine LOCC-based witness arguments in realistic bound-matter platforms.
1 Introduction
The possibility that gravity can generate entanglement between two otherwise isolated quantum systems has been widely discussed as a potential witness of the nonclassicality of the gravitational interaction. In the standard narrative, if two systems $A$ and $B$ start in a separable state and interact only via a mediator that transmits classical information (LOCC-type mediation), then entanglement cannot be created. This motivates proposals in which two macroscopic masses are placed in spatial superpositions and their mutual gravitational interaction is inferred through entanglement generation.
Aziz and Howl (hereafter AH) challenge the universality of this inference by arguing that, when matter is treated as a scalar field in a perturbative QFT framework, a classical gravitational field can still induce entanglement through higher-order processes involving virtual matter exchanging between the two systems [aziz2025classical]. On the other hand, several recent comments have questioned whether the specific nonrelativistic limit considered in AH actually generates entanglement, and emphasized that any entangling capability—if present—would originate from a quantum matter channel rather than from a purely classical mediator [marletto2025classicalgravitymediateentanglement, diósi2025noclassicalgravitydoes, sienicki2025commentclassicalgravityquantummatterclaimsgravitymediated].
We agree that such “virtual-matter” terms do appear in perturbation theory in a classical gravitational background and, at least formally, can contribute to entanglement generation. However, the relevant channel is not intrinsically a QFT phenomenon, but is essentially tunneling/evanescent propagation in quantum mechanics. In realistic macroscopic objects, microscopic constituents are bound and localized. Once binding is included, any “virtual-matter” channel across a macroscopic separation is exponentially suppressed, so the corresponding entanglement mechanism is operationally irrelevant in the parameter regimes of gravitational-entanglement proposals. Moreover, this non-LOCC matter-exchange channel can generate entanglement even without gravity, so it should not be attributed to “classical gravity” mediation.
2 QFT is not necessary
AH exhibit entanglement generation by evaluating Feynman diagrams in perturbative QFT. While QFT is indispensable in many contexts—notably when particle number can fluctuate and when scattering processes admit a perturbative expansion—the specific mechanism highlighted by AH can be naturally understood in simpler quantum-mechanical terms.
In particular, the diagram responsible for the classical-gravity-induced entanglement in AH (the upper-right diagram in Fig. 2 of [aziz2025classical], reproduced schematically in our Fig. 1) involves no particle creation or annihilation. Its physical content is a coherent matter-mediated amplitude connecting two spatial regions, which in quantum mechanics is precisely an evanescent/tunneling process. Put differently, what is called a “virtual particle” in a diagrammatic expansion corresponds, in a first-quantized description, to propagation through off-shell intermediate states (tunneling) rather than to an essentially field-theoretic effect.
A more serious issue concerns the microscopic modeling of matter. AH treat matter as a massive scalar field $\phi$ that propagates nearly freely in space (their Eq. (3)). Such an approximation may be natural for elementary-particle scattering, but it is not an appropriate model for macroscopic solids. To bind atoms into a solid, one must include strong interactions or, at minimum, an effective binding/localizing potential. Without such binding, the microscopic wavefunctions would delocalize, and one is no longer describing a solid object.
Figure 1: This figure describes AH’s channel generating quantum entanglement between two objects. In the middle of this figure is the Feynman diagram of particle tunneling (Fig. 2 in AH’s paper). They assume all particles are free, but in a solid, atoms can propagate only if they are excited by a virtual energy $E_{b}$, which leads to an exponential-decay factor.
3 Binding implies exponential suppression
In standard settings, tunneling/evanescent propagation over large distances is exponentially suppressed. This suppression is absent in AH’s process because binding/localization is not included in the microscopic model. Realistic matters are condensed many-body systems: microscopic constituents are localized by binding and lattice potentials with a large internal energy scale. In such systems, coherent propagation between separated bodies is evanescent and short-ranged.
A minimal estimate follows from the WKB approximation. If a constituent of mass $m$ must access a virtual energy of order $E_{b}$ to leave its bound region, the evanescent wavevector is
$\kappa\sim\frac{\sqrt{2mE_{b}}}{\hbar},$ (1)
so the amplitude to connect two bodies separated by $d$ scales as
$\mathcal{A}(d)\propto e^{-\kappa d}=e^{-d/\ell},\qquad\ell\equiv\frac{\hbar}{\sqrt{2mE_{b}}}.$ (2)
For typical atom mass $m\sim 10^{-27}~{\rm kg}$ and binding energy $E_{b}\sim 1~{\rm eV}$, one finds $\ell$ at the picometer scale (e.g. $\ell\sim 10^{-11}~{\rm m}$), so $e^{-d/\ell}$ is negligible for macroscopic separations.
The quantum field–theoretic (QFT) description is consistent with this picture. A finite binding energy $E_{b}$, which raises the potential outside the masses, modifies the particle propagator across the bodies. This escaping energy can be interpreted as an effective mass shift in the propagator denominator,
$m^{\prime}c^{2}=mc^{2}+E_{b}.$ (3)
and we will consider it within the non-relativistic regime $E_{b}\ll mc^{2}$.
Because the mass shell itself is not shifted by $E_{b}$, this modification produces a mismatch in the exponential factors appearing in the relevant Feynman diagrams, which is central to AH’s analysis of distance-dependent decay. In their treatment, exponential factors arising from the contraction of the initial and final states cancel exactly against those of the propagator. Here, this cancellation fails, leaving a residual factor $e^{-\sqrt{m^{\prime 2}-m^{2}}\,\frac{c}{\hbar}d}\approx e^{-\sqrt{2mE_{b}}\,\frac{d}{\hbar}}$, which reproduces the suppression identified in the quantum-mechanical regime.
One can compare this result to Eq. (69) of AH’s Supplementary Material of [aziz2025classical], where our pole is fixed at $k_{0}=m^{\prime}c/\hbar$ rather than $mc/\hbar$. The absence of this decay in AH’s Eq. (69) arises from neglecting the role of the binding force in suppressing long-range propagation, treating it solely as a constraint on the initial and final states (Eq. (9) in the Supplementary Material). Incorporating a realistic $E_{b}$ restores an evanescent factor and renders the effective matter correlator short-ranged, scaling as $\sim e^{-d/\ell}$. Consequently, any matter-mediated entangling contribution requiring coherent propagation between macroscopic solids is exponentially suppressed in realistic platforms.
4 Quantum tunneling is not LOCC
This also clarifies why entanglement arises in AH’s analysis. The Feynman diagram Fig. 1 means that two objects can exchange particles. When particles are described by quantum mechanics, exchanging particles is beyond classical communication. We give a simple example in their two-object system. Suppose the first object has two energy levels $1L,1R$, and the second object has two energy levels $2L,2R$. Let the initial state be $|N_{1L}=N,N_{1R}=0\rangle\otimes|N_{2L}=N,N_{2R}=0\rangle$, and after a certain period of time, there is a small tunneling amplitude $\alpha\ll 1$ that an atom from $1L$ tunneling to $2R$. Then, up to normalization, the state becomes
$|\Psi\rangle\propto|N,0\rangle\otimes|N,0\rangle+\alpha\,|N-1,0\rangle\otimes|N,1\rangle,$ (4)
which is clearly entangled. This channel is essentially the same as AH’s virtual-matter channel, while it is simpler because no interaction is assumed. Even if there is no gravity, tunneling itself is able to generate entanglement. Nonetheless, this effect is exponentially suppressed in real experiments as long as the objects are solid or bounded.
Acknowledgements. We acknowledge the support of National Natural Science Foundation of China (No. 12374325), Young Elite Scientists Sponsorship Program by CAST (Grant No. 2023QNRC001), and Beijing Municipal Science and Technology Commission (Grant No. Z221100002722011). Z. Han acknowledges the support of the National Natural Science Foundation of China (Grant No. T2225008) and the Tsinghua University Dushi Program.